OpenStudy (anonymous):

A box has 1 red marble, 1 blue marble, and 4 green marbles of the same shape and size. Dan draws a marble randomly from the box, replaces it, and then draws another marble randomly. What is the probability of drawing two green marbles in a row? Explain your answer.

OpenStudy (anonymous):

Wouldnt It be 4/6 * 4/6=16/36=4/9

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct